Simplify and solve the following linear equations. 
3(5z – 7) – 2(9z – 11) = 4(8z – 13) – 17 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3(5z–7)–2(9z–11)=4(8z–13)–17
⇒15z-21-18z+22=32z-52-17
⇒15z-18z-32z=-52-17+21-22
⇒-35z=-70
⇒z = -70/-35
⇒z=2

Thanks